OpenStudy (anonymous):

(x - a) (x + a) = x2 – a2 find 62 – 52.

OpenStudy (zehanz):

You probably mean:\[(x-a)(x+a)=x^2-a^2\]and:\[6^2-5^2\] Now this is a little far-fetched, because everybody knows that 6²-5²=36-25=11, but just for fun:\[6^2-5^2=(6-5)(6+5)=...\]You can see where this is heading to, I guess...